OpenStudy (anonymous):

vertex, focus, directrix, and focal width: x=4y^2?

OpenStudy (jdoe0001):

I notice that the "y" variable is the squared component, so for that, a "focus form" equation will be like => \(\bf (y-k)^2=4p(x-h)\) so, simplify \(\bf x=4y^2\) to \(\bf (y-k)^2=4p(x-h)\)

OpenStudy (jdoe0001):

keep in mind that in \(\bf (y-k)^2=4p(x-h)\) (h, k) = vertex coordintes p = distance from the vertex to the focus the distance from the focus to the vertex, is the same distance as from the vertex to the directrix in the opposite direction

OpenStudy (anonymous):

yeah I get what each thing means I just cant figure out how to find the focus, vertex, directrix and focal width, I have to actually see it done to understand how to do it.

OpenStudy (anonymous):

If you could guide me or something that'd be great.